OpenStudy (anonymous):

121 22' 36" is the same as____ (round to the nearest hundredth of a degree)

jimthompson5910 (jim_thompson5910):

Similar Example (just with different numbers) \[\Large 247^{\circ} 18' 42'' = 247^{\circ} \color{red}{18}' \color{blue}{42}''\] \[\Large 247^{\circ} 18' 42'' = 247+\frac{\color{red}{18}}{60} +\frac{\color{blue}{42}}{3600}\] \[\Large 247^{\circ} 18' 42'' \approx 247+0.3+0.011667\] \[\Large 247^{\circ} 18' 42'' \approx 247.311667^{\circ}\]

jimthompson5910 (jim_thompson5910):

The 60 under the 18 refers to the fact that there are 60 minutes in a degree The 3600 is from the fact that there are 60*60 = 3600 seconds in a degree